If the labor force of 160 million people is growing by 1.2 percent this year, how many new jobs have to be created each month to keep unemployment from increasing?

Instructions: Enter your response as a whole number.

New jobs per month:  
Answer #1

Labor force = 160 million = 160 , 000 ,000

Increase by 1.2 %

Increase after 1 year = 160,000,000 * 1.2/100 = 1,920,000

New jobs per month = 1,920,000 / 12 = 160,000